At 20 degree celsius two balloons of equal volume and porosity are filled to pressure of 2 atm. one with 14kg Nitrogen and other with 1kg hydrogen. the nitrogen molecule leaks to a pressure of 1/2 atm in 1 hr. How long will it take for hydrogen balloon to reach a pressure of 1/2 atm

Grade:11

2 Answers

Carl Jhonson
7 years ago
The ans is 16minI dont know how to solve it but it can be done by grahms law of effusion and diffusion . r/R=(m/M)½
Nayana
6 years ago
NO.of moles of N2=500No.of moles of H2=500Find the initial volume using any of the given data as no.of moles,pressure, temperature are initially the same.PV=notV=6013.83LNow considering the temperature to remain constant ,the volume changes as pressure doesP1×V1=P2×V22×6013.83=0.5×V2V2=24055.3Lchange in volume=24055.3-6013.83 =18041.47LThis change in volume remains same for both N2 and H2 as final pressures are same.Time for N2=60 min (given )Rate of N2 (r)=change in volume/time =300.7L/minApply graham`s law300.7/R=(2÷28)^0.5R (rate of H2)=1125.12L/minAs mentioned above change in volume remains sameR=change in volume/time1125.12=18041.47/timeTime=16.035~16min